OpenStudy (anonymous):

Choose the correct answer. Yves bought 420 tropical fish for a museum display. He bought 6 times as many parrotfish as angelfish. How many of each type of fish did he buy? Which is a system of equations to model the problem if x represents the number of angelfish Yves bought and y represents the number of parrotfish he bought? A. x + y = 6 y = 420x B. x + 6y = 420 y = 6x C. x – y = 420 y = 6x D. x + y = 420 y = 6x @Directrix

Directrix (directrix):

You'll have to use the same assignment of variables that "they" did x represents the number of angelfish y represents the number of parrotfish

Directrix (directrix):

So, x + y + ?

Directrix (directrix):

Re-read the problem: Yves bought 420 tropical fish for a museum display.

OpenStudy (anonymous):

i think it is B

Directrix (directrix):

We are not doing answers yet. We are doing work.

Directrix (directrix):

So, x + y = ?

OpenStudy (anonymous):

420+6=426

OpenStudy (anonymous):

i don't really get it um hold up

OpenStudy (anonymous):

x+y=420

Directrix (directrix):

How many fish did this dude buy. Look at the question.

OpenStudy (anonymous):

he bought 420

Directrix (directrix):

Again, >.Yves bought 420 tropical fish for a museum display.

OpenStudy (anonymous):

he bought 420 tropcal fish

OpenStudy (anonymous):

ok ik that i just get confused on the rest

OpenStudy (anonymous):

am i suppose to multiply 420x6?

Directrix (directrix):

x+y=420 is the first equation. Now, for the second.

Directrix (directrix):

bought 6 times as many parrotfish as angelfish. Parrotfish = 6*Angelfish

OpenStudy (anonymous):

so D

Directrix (directrix):

We are not doing answers yet. We are getting the second equation.

Directrix (directrix):

x represents the number of angelfish y represents the number of parrotfish Parrotfish = 6*Angelfish

Directrix (directrix):

y = ?

OpenStudy (anonymous):

so it would be y=6

OpenStudy (anonymous):

so it would be y=6x

Directrix (directrix):

No, that would mean the dude bought 6 parrotfish. ------------------------------------- He bought 6 times the number of angelfish.

Directrix (directrix):

So, y = 6 * ?

OpenStudy (anonymous):

x represents the number of angelfish y represents the number of parrotfish Parrotfish = 6*Angelfish as you said

Directrix (directrix):

What is the variable for angelfish

OpenStudy (anonymous):

so we would replace the parrotfish with y so we get y=6 then we replace the angelfish with x so we would get y=6x

Directrix (directrix):

Correct. So, which option is that?

OpenStudy (anonymous):

D

Directrix (directrix):

That is what I think.

OpenStudy (anonymous):

thanks

Directrix (directrix):

You are welcome.

OpenStudy (anonymous):

um the next question is kind of like this but a bit different

Directrix (directrix):

Same song, different verse.

OpenStudy (anonymous):

hehe

Directrix (directrix):

Close and post.

OpenStudy (anonymous):

that's really funny